OpenStudy (anonymous):

Use the properties of logarithms to expand the expression ln z(z-1)^2, z>1?

OpenStudy (anonymous):

Check it on yahoo . It says " = Ln (z) + Ln (z-1)^2 = Ln (z) + 2 Ln (z-1) "

OpenStudy (anonymous):

that's the question I posted

OpenStudy (anonymous):

and no one would explain

OpenStudy (anonymous):

amofide xD

OpenStudy (amorfide):

you have rules for logs if you have lnx² that is the same as saying 2lnx you can bring the power to the front of the log

OpenStudy (anonymous):

but how do you separate it? Like Ln (z) + Ln (z-1)^2

OpenStudy (anonymous):

i mean Ln (z) + 2 Ln (z-1)

OpenStudy (amorfide):

also you started with ln z(z-1)^2 there is another rule for logs if you have lna + ln b this is equal to lnab so you have ln z(z-1)^2 this is being multiplied, so we know we have two logs added together lnz + ln (z-1)²

OpenStudy (anonymous):

Ooooooooooooooh

OpenStudy (anonymous):

Thanks!

OpenStudy (amorfide):

anytime :D so lna+lnb=lnab lna-lnb=lna/b lna^b=blna i think there is another one i forget...